OpenStudy (anonymous):

You are given P(A∪B)=0.7 and P(A∪B' )=0.9. Determine P (A). A) 0.6 B) 0.3 C) 0.4 D) 0.2 E) 0.8

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

P((A∪B)∪(A∪B' ) ) =P(A∪B)+P(A∪B' )-P((A∪B)∩(A∪B' ) ) thus P(U)=P(A∪B)+P(A∪B' )-P(A), where P(U) is the probability of the universal set. 1=0.7+0.9-P(A) finally P(A)=0.6